OpenStudy (anonymous):

Is it true that 2 x 2 = 5 ??? (I Can Prove It!)?

OpenStudy (anonymous):

Omg I can't wait you must be a genius

OpenStudy (osuxbucks15):

yes

OpenStudy (osuxbucks15):

now prove it

OpenStudy (anonymous):

WHy ,,

Parth (parthkohli):

You can prove that \(2 \times 2 \ne 5\) by a contradiction.

Parth (parthkohli):

First assume that \(2 \times 2 = 5\)

Parth (parthkohli):

Guess what you can do here ;)

OpenStudy (anonymous):

hahahahha go Parth!

OpenStudy (anonymous):

the answer is sort of a joke!!

OpenStudy (anonymous):

WHy 2x2 is = 5??

Parth (parthkohli):

2 * 2 = 2 + 2 = 5

Parth (parthkohli):